# High-Resolution Flood Probability Mapping Using Generative Machine Learning with Large-Scale Synthetic Precipitation and Inundation Data

### v1

@@ 2.4 Rainfall Data Augmentation via CTGAN
The rainfall events considered in simulating flood scenarios with the physics-based model were limited to 592 events. To increase the number of rainfall events and generate synthetic rainfall events in our study, we employed Conditional Generative Adversarial Networks (CTGAN) from SDV [22, 23] for tabular data augmentation. CTGAN extends the GAN framework to handle structured data, particularly tabular datasets, making it highly effective for generating synthetic data in scenarios with small, imbalanced datasets. One of CTGAN’s strengths is its ability to support the specification of constraints and conditions during the data generation process, enabling the preservation of specific relationships or attributes present in the original dataset.
We converted the true cell-level data into point-level data using centroids to facilitate generation. Consequently, the input attributes for CTGAN included latitude, longitude, cumulative precipitation, peak precipitation, and duration. To enhance the quality of the synthetic records, we introduced several constraints to ensure the synthetic point data adhered to the following conditions: (1) the data remained within specified geographical boundaries, (2) the cumulative precipitation exceeded the peak precipitation, and (3) the peak precipitation surpassed the unit precipitation, calculated by dividing the cumulative precipitation by the duration.
To determine the optimal hyperparameters for the generator and discriminator learning rates, as well as the stoppage epochs, we conducted a comprehensive hyperparameter grid search. The grid search was crucial due to the complexity of balancing the learning rates for the generator and discriminator, ensuring both networks improved steadily without overpowering each other. Additionally, selecting appropriate stoppage epochs enabled us to terminate the training process at the most accurate point, preventing overfitting or underfitting. The optimal checkpoint was selected based on the corresponding synthetic dataset achieving the highest average marginal distribution, as evaluated using the Kolmogorov-Smirnov statistic (KS statistic) [24]. The KS statistic $K$ for a particular feature $x_{i}$ between the training dataset and the synthetic dataset is defined as follows:
@@
Where $N$ is the number of features. A higher score indicates a higher quality of the synthetic data, signifying a closer approximation to the true data distribution. The outcome of this step is a comprehensive synthetic point cloud containing 10,000,000 synthetic rainfall records generated by CTGAN, representing various precipitation durations. This diverse point cloud presents a significant challenge in processing and formulating individual precipitation events under varying conditions.
@@ 2.5 Rainfall Events Generation
In the synthetic point dataset, which comprises collections of precipitation records, we have predefined thresholds for cumulative precipitation, peak precipitation, and duration—categorized as Low, Medium, and High (LMH)—for each cell mesh. These categories are established based on the following conditions:
@@
Where $\mu_{i}$ is the mean of a precipitation-based feature $x_{i}$ from training dataset, and $\sigma_{i}$ refers to its standard deviation. $\theta_{1}$ and $\theta_{2}$ are two positive constants. Consequently, for each cell mesh, we established low (L), medium (M), and high (H) thresholds based on the distribution of precipitation-based features (cumulative precipitation, peak precipitation and duration) from the training dataset. These thresholds were then applied to the synthetic point dataset generated as described in section 2.4. We conducted a comprehensive iteration over all synthetic point records, retaining the index of a point record to corresponding cell-level feature pool only if all its synthetic precipitation-based features consistently matched within the same LMH classification. This rigorous iterative screening ensured that each cell’s pool included indices of synthetic records that not only covered a range of precipitation levels but also met stringent quality criteria.
@@
With the establishment of the synthetic indices pool for each cell, we were enabled to generate global precipitation events across the study area. For each event generation, a specific duration was sampled randomly from the aggregated pool to set as the global rainfall duration, constraining the potential pool. Subsequently, each cell randomly sampled a synthetic record from its tailored pool to represent the entire cell mesh. In instances where some cells returned an empty record, we employed a K-Nearest Neighbors (KNN) approach [25] to generate records by averaging the attributes of surrounding neighbors, effectively filling any gaps. This strategic sampling allowed us to create isolated synthetic precipitation events from a large point cloud, transitioning from point-level to cell-level representation. To enhance the approximation of true precipitation events, we applied the KNN method as a smoother [26] to both cumulative and peak precipitation data. Additionally, we computed heavy precipitation ratios from these smoothed values to highlight interactions among various watersheds. Following this cell-level processing, we utilized the depth estimator with optimal predictive performance as identified in Section 2.3 to generate inundation depths from the aggregated synthetic features, thus completing the precipitation-flood event. After thousands of iterations of individual rainfall event generation, we built synthetic inundation depth distribution for each cell, enabling the estimation of the frequency of rainfall events that cause certain flood depth levels in different areas in producing the synthetic flood probability maps.
@@
To rigorously assess the event-generation performance, we evaluated the statistical similarity between the true inundation depth distribution and the synthetic inundation depth distribution. We implemented interpolation using the formula in Eq. 5 for the synthetic depth distribution $\mathbf{d_{s}}$ to ensure the same dimension as the training depth distribution $\mathbf{d_{t}}$, facilitating the evaluation. Let $\mathbf{d_{t}}=[d_{t}(1),d_{t}(2),\ldots,d_{t}(m)]$ be the training depth vector with $m$ records, and $\mathbf{d_{s}}=[d_{s}(1),d_{s}(2),\ldots,d_{s}(n)]$ be the synthetic depth vector with $n$ synthetic records. We define an interpolation function $f$ such that:
@@
This process ensures that $\mathbf{d_{t}}$ and $\mathbf{d_{s}^{\prime}}$ are of the same length, enabling valid comparisons using multiple statistical measures. Cosine similarity [27] assesses their directional similarity, Pearson correlation [28] evaluates the linear relationship between the two distributions, and Kullback-Leibler (KL) divergence [29] quantifies how one probability distribution diverges from the other. The corresponding formula is shown below:
@@
$\text{Correlation}(\mathbf{d_{t}},\mathbf{d_{s}^{\prime}})=\frac{\sum_{i=1}^{m}(d_{t}(i)-\bar{d_{t}})(d_{s}^{\prime}(i)-\bar{d_{s}^{\prime}})}{\sqrt{\sum_{i=1}^{m}(d_{t}(i)-\bar{d_{t}})^{2}\sum_{i=1}^{m}(d_{s}^{\prime}(i)-\bar{d_{s}^{\prime}})^{2}}}$ (7)
$D_{KL}(\mathbf{p}\parallel\mathbf{q})=\sum_{i=1}^{m}p(i)\log\frac{p(i)}{q(i)},\ p(i)=\frac{d_{t}(i)}{\sum^{m}_{k=1}d_{t}(k)},\ q(i)=\frac{d_{s}^{\prime}(i)}{\sum^{m}_{k=1}d_{s}^{\prime}(k)}$ (8)
where $\mathbf{p}$ and $\mathbf{q}$ are the normalized versions of $\mathbf{d_{t}}$ and $\mathbf{d_{s}^{\prime}}$, respectively. Besides the averaged synthetic depth distribution across the study region, we further explore the performance by aggregating the results into channel and non-channel levels, as described in Section 2.3. This detailed analysis allows us to assess how well the synthetic data captures the unique characteristics of different regions in the study area.
@@ 3.1 Selection of Depth Estimator
To understand the advancements of tree-based models and transformers in regression problems, specifically in the context of inundation depth estimation, we compared the predictive performance of two universal models, XGBoost Regressor and Regression Transformer, with one cell-wise model, MaxFloodCast V2. MaxFloodCast V2 utilizes an XGBoost-based architecture and incorporates heavy precipitation ratio features. We trained the Regression Transformer, configured with 4 layers in the Transformer encoder, a dimensionality of 128, 2,048 in the feed-forward layer, and 8 attention heads with 25% dropout, on 8 NVIDIA RTX A100 GPUs. The XGBoost Regressor was trained on an NVIDIA RTX A6000, while MaxFloodCast V2 was trained in parallel on an AMD EPYC 7702P 64-Core Processor, sharing the same hyperparameter settings as the XGBoost Regressor. The objective function minimized squared errors with a learning rate of 0.01. The XGBoost model built 1000 trees with a maximum depth of 5, incorporating L1 regularization to prevent overfitting and a subsample ratio of 0.3 to introduce non-linearity and enhance robustness. Following the sampling strategy mentioned in Section 2.2, we configured 252 true precipitation-flood events for training, 63 events for validation, and the remaining events as the test set. The comparison results are presented in Table 3, which assesses the overall, channel, and non-channel RMSE, as well as the $R^{2}$ score.
Table 3: Depth Estimator Comparison on true test dataset.
Model Level Overal RMSE Overal $R2$ Channel RMSE Channel $R^{2}$ Non-channel RMSE Non-channel $R^{2}$
Regression Transformer[18] Universal 2.6634 0.6537 4.1264 0.7278 2.4133 0.4928
XGBoost Regressor[16] Universal 2.6916 0.6463 4.2755 0.7078 2.4153 0.4920
MaxFloodCast V2[20] Cell-wise 0.6996 0.9189 1.7280 0.9107 0.5680 0.9200
@@
The resulted comparison reveals significant differences in the predictive performance between the universal models and the cell-wise model. The Regression Transformer and XGBoost Regressor both show substantial errors, with overall RMSE scores above 2 ft, indicating they struggle to accurately predict inundation depths. Their performance is particularly poor in channel cells, where RMSE values exceed 4 ft. In contrast, MaxFloodCast V2 demonstrates superior performance with an overall RMSE of 0.6996 ft and an $R^{2}$ value of 0.9189, significantly outperforming the universal models. It achieves lower RMSE in both channel (1.7280 ft) and non-channel cells (0.5680 ft), highlighting its accuracy and reliability. The cell-wise approach of MaxFloodCast V2, which tailors the model to the specific characteristics of each cell, allows it to better capture local variations in precipitation and terrain features, resulting in more accurate predictions. Also, the incorporation of heavy precipitation ratio features enhances its ability to represent interactions between watersheds. The dataset used for training, generated by the physics-based model, is both limited and imbalanced, comprising 90 true events. This limitation affects the universal models more significantly, as they are less adept at handling such imbalances compared to the specialized cell-wise approach of MaxFloodCast V2. Overall, the results highlight the effectiveness of the cell-wise approach in flood depth estimation, particularly in regions with complex hydrological dynamics, making MaxFloodCast V2 the optimal depth estimator for the Flood-Precip GAN model.
@@
Figure 3: Grid search results for CTGAN hyperparameters. The optimal hyperparameter set was selected based on the highest average marginal distribution scores across all synthetic precipitation-based features generated by the best-performing CTGAN configurations. The key hyperparameters explored in the grid search include the generator and discriminator learning rates, as well as the number of training epochs.
@@ 3.2 CTGAN-generated Precipitation Records Quality Review
We strategically sampled 90 true events following the ratio described in Section 2.2 and configured 36 hyperparameter sets for our grid search. This involved combining six pairs of learning rate settings with six distinct stoppage epochs, increasing from 50 epochs to 300 epochs in increments of 50, and modeling on NVIDIA’s RTX A6000 GPU. This meticulous grid search focused on generating records under the condition $\text{duration}\geq 1$, targeting the potential simulation of global precipitation events. As shown in Fig. 3, we identified the optimal checkpoint with a generator learning rate of $10^{-2}$, a discriminator learning rate of $10^{-3}$, and early stopping at 250 epochs. Notably, the generator learning rate is set to be 10 times higher than that of the discriminator to prevent the generator from being overpowered in the early stages before it can effectively learn the underlying data patterns. Ultimately, both the generator and discriminator converged before reaching the 300-epoch limit. The optimal settings enabled a direct comparison of cumulative precipitation, peak precipitation, and duration between the training dataset and the 10,000,000 synthetic data points generated by CTGAN, as depicted in Fig. 4. This configuration produced an average marginal distribution score of 0.802, demonstrating a high fidelity of the synthetic data, particularly for cumulative and peak precipitation when compared to the original dataset. As shown in Fig. 4(c), we enforced a constraint in CTGAN that limits $\text{duration}\leq 1$, since our focus is primarily on global precipitation event generation. Additionally, the CTGAN model maintains the flexibility to generate local precipitation events by allowing cumulative precipitation to drop to zero. Moving forward, the challenge remains to construct isolated precipitation-depth events from the synthetic point dataset that closely correlate with true events, thus advancing the goals of this research in generating probabilistic flood maps.
Figure 4: Stacked Marginal Distribution Comparison of Three Synthetic Features between training data and synthetic data: (a) cumulative precipitation, (b) peak precipitation and (c) duration. The distribution of the training dataset is represented in gray, while the synthetic distribution is depicted in light blue. Distributions were generated using the Synthetic Data Vault (SDV) [22].

### v2

@@ 2.4 Rainfall Data Augmentation via CTGAN
Given the limited physics-based model simulated precipitation-flood events, to generate relatively robust synthetic rainfall events in our study, we strategically sampled 315 events as explained in Section 2.2 and then employed Conditional Generative Adversarial Networks (CTGAN) from SDV 21, 32 for tabular data augmentation. CTGAN extends the GAN framework to handle structured data, particularly tabular datasets, making it highly effective for generating synthetic data in scenarios with small, imbalanced datasets. One of CTGAN’s strengths is its ability to support the specification of constraints and conditions during the data generation process, enabling the preservation of specific relationships or attributes present in the original dataset.
We converted the training cell-level data into point-level data using centroids to facilitate generation. Consequently, the input attributes for CTGAN included latitude, longitude, cumulative precipitation, peak precipitation, and precipitation duration. To enhance the quality of the synthetic records, we constrained conditional sampling to ensure that synthetic points satisfied the following conditions: (1) sampled points are inside the study area (Fig. 2); (2) for a sampled point the cumulative precipitation is always greater than the peak precipitation value; (3) for a sampled point the peak precipitation is greater than the average precipitation intensity, calculated by dividing the cumulative precipitation by the event duration.
To determine the optimal hyperparameters for the generator and discriminator learning rates, as well as the stoppage epochs, we conducted a comprehensive hyperparameter grid search to explore the optimal learning rates for the generator and discriminator, ensuring both networks improved steadily without overpowering each other. Additionally, selecting appropriate stoppage epochs enabled us to terminate the training process at the most accurate point, preventing overfitting or underfitting. The optimal checkpoint was selected based on the corresponding synthetic dataset achieving the highest average marginal distribution, as evaluated using the Kolmogorov-Smirnov statistic (KS statistic) 18. The KS statistic $K$ for a particular feature $x_{i}$ between the training dataset and the synthetic dataset is defined as follows:
@@
Where $N$ is the number of features. A higher score indicates a higher quality of the synthetic data, signifying a closer approximation to the training data distribution. The outcome of this step is a comprehensive synthetic point cloud containing 10,000,000 synthetic rainfall records generated by CTGAN, representing various precipitation durations. This diverse point cloud presents a significant challenge in processing and formulating individual precipitation events under varying conditions.
@@ 2.5 Rainfall Events Generation
In the synthetic point dataset, which comprises collections of precipitation records, we have predefined thresholds for cumulative precipitation, peak intensity, and duration respectively—categorized as Low, Medium, and High (LMH)—for each cell mesh based on training data. These categories are established based on the conditions shown on table 2.
@@
Where $\mu_{i}$ is the mean of a precipitation-based feature $x_{i}$ from training dataset, and $\sigma_{i}$ refers to its standard deviation. $\theta_{1}$ and $\theta_{2}$ are two positive constants. Consequently, for each cell mesh, we defined low (L), medium (M), and high (H) thresholds based on the distribution of precipitation-related features—cumulative precipitation, peak precipitation, and duration—derived from the training dataset. These thresholds were then applied to the synthetic point dataset generated as described in Section 2.4. For each cell, we constructed a corresponding precipitation feature-level pool comprising all possible LMH combinations of cumulative precipitation, peak precipitation, and duration, resulting in a total of 27 combinations (i.e., LLL, LLM, …, HHH.). To ensure consistency and representativeness, we systematically iterated over all synthetic point records, retaining indices of synthetic points that fell within the cell boundary while preserving the original distribution of the pool.
@@
Establishing the synthetic indices pool for each cell enabled to generate global precipitation events with various precipitation distributions across the study area. To generate an event a specific duration was sampled randomly from the aggregated pool and set as the global rainfall duration, reducing the possible cumulative-peak-duration combinations from 27 to 9 (e.g., if the input synthetic duration is H, then the potential set of combination is LLH, LMH, LHH, MLH, MMH, MHH, HLH, HMH, HHH). Subsequently, a synthetic record was randomly sampled for each cell from its own pool. The sampling was repeated for all the cells of the mesh. This strategic sampling allowed to create single synthetic precipitation events from a large point cloud, transitioning from point-level to cell-level representation in accord with the original distribution of precipitation characteristics in the study area. To enhance the approximation of training precipitation events, we applied the KNN method as a smoother 31 to cumulative and peak precipitation features respectively. Additionally, we computed HCPR and HPPR from these smoothed values. Following this cell-level processing, we utilized the depth estimator with optimal predictive performance as identified in Section 2.3 to generate inundation depths resulting from each synthetic precipitation event. The depth estimator was applied to the thousands of generated rainfall events and produced a synthetic inundation depth distribution in each cell.
enabling the estimation of the frequency of rainfall events that cause certain flood depth levels in different areas in producing the synthetic flood probability maps.
@@
To assess the event-generation performance, we evaluated the statistical similarity between the inundation depth distribution of the training set of rainfall events and the inundation depth distribution of the synthetic set. We implemented interpolation using the formula in Eq. 5 for the synthetic set depth distribution $\mathbf{d_{s}}$ to ensure the same dimension as the training set depth distribution $\mathbf{d_{t}}$, facilitating the evaluation. Let $\mathbf{d_{t}}=[d_{t}(1),d_{t}(2),\ldots,d_{t}(m)]$ be the training set depth vector with $m$ records, and $\mathbf{d_{s}}=[d_{s}(1),d_{s}(2),\ldots,d_{s}(n)]$ be the synthetic set depth vector with $n$ synthetic records. We define an interpolation function $f$ such that:
@@
This process ensures that $\mathbf{d_{t}}$ and $\mathbf{d_{s}^{\prime}}$ are of the same length, enabling valid comparisons using multiple statistical measures. Cosine similarity 30 assesses their directional similarity. Pearson correlation 7 evaluates the linear relationship between the two distributions. Finally, Kullback-Leibler (KL) divergence 11 quantifies how one probability distribution diverges from the other. The corresponding formulas can be written as:
@@
$\text{Corr}(\mathbf{d_{t}},\mathbf{d_{s}^{\prime}})=\frac{\sum_{i=1}^{m}(d_{t}(i)-\bar{d_{t}})(d_{s}^{\prime}(i)-\bar{d_{s}^{\prime}})}{\sqrt{\sum_{i=1}^{m}(d_{t}(i)-\bar{d_{t}})^{2}\sum_{i=1}^{m}(d_{s}^{\prime}(i)-\bar{d_{s}^{\prime}})^{2}}}$ (7)
$D_{KL}(\mathbf{p}\parallel\mathbf{q})$ $=\sum_{i=1}^{m}p(i)\log\frac{p(i)}{q(i)},$ (8)
$p(i)$ $=\frac{d_{t}(i)}{\sum^{m}_{k=1}d_{t}(k)},$ (9)
$q(i)$ $=\frac{d_{s}^{\prime}(i)}{\sum^{m}_{k=1}d_{s}^{\prime}(k)}$ (10)
where $\mathbf{p}$ and $\mathbf{q}$ are the normalized versions of $\mathbf{d_{t}}$ and $\mathbf{d_{s}^{\prime}}$, respectively, as represented in Eqs. 9 and 10. Besides the averaged synthetic depth distribution across the study region, we further explore the performance by aggregating the results into channel and non-channel levels, as described in Section 2.3. This detailed analysis allows to assess how well the synthetic data captures the unique characteristics of different regions in the study area.
@@ 3.1 Selection of Depth Estimator
To understand the advancements of tree-based models and transformers in regression problems, specifically in the context of inundation depth estimation, we compared the predictive performance of two universal models, XGBoost Regressor and Regression Transformer, with one cell-wise model, MaxFloodCast V2. MaxFloodCast V2 utilizes an XGBoost-based architecture and incorporates heavy precipitation ratio features. We trained the Regression Transformer using an encoder with 4 layers, a model dimensionality of 128, a feed-forward layer size of 2,048, and 8 attention heads with a 25% dropout rate on 8 NVIDIA RTX A100 GPUs. The XGBoost Regressor was trained on an NVIDIA RTX A6000, while MaxFloodCast V2 was trained in parallel on an AMD EPYC 7702P 64-Core Processor, sharing the same hyperparameter settings as the XGBoost Regressor. The model’s objective function aimed to minimize RMSE, utilizing a learning rate of 0.01. The XGBoost algorithm generated 1000 trees with a maximum depth of 5, applying L1 regularization to mitigate overfitting and employing a subsample ratio of 0.3 to introduce non-linearity and improve model robustness. Following the sampling strategy described in Section 2.2, we configured 252 precipitation-flood events for training, 63 events for validation, and the remaining events as the test set. The comparison results are presented in Table 3, which assesses the overall, channel, and non-channel RMSE, as well as the $R^{2}$ score.
Table 3: Depth Estimator Comparison on test dataset.
Model Level Overal RMSE Overal $R^{2}$ Channel RMSE Channel $R^{2}$ Non-channel RMSE Non-channel $R^{2}$ Time
Regression Transformer 29 Universal 2.6634 0.6537 4.1264 0.7278 2.4133 0.4928 5.8s
XGBoost Regressor 3 Universal 2.6916 0.6463 4.2755 0.7078 2.4153 0.4920 4.1s
MaxFloodCast V2 15 Cell-wise 0.6996 0.9189 1.7280 0.9107 0.5680 0.9200 10.6s
@@
The comparison of results reveals significant differences in the predictive performance between the universal models and the cell-wise model. The Regression Transformer and XGBoost Regressor both show substantial errors, with overall RMSE scores above 2 ft, indicating they struggle to accurately predict inundation depths. Their performance is particularly poor in channel cells, where RMSE values exceed 4 ft. In contrast, MaxFloodCast V2 demonstrates superior performance with an overall RMSE of 0.6996 ft and an $R^{2}$ value of 0.9189, significantly outperforming the universal models. It achieves lower RMSE in both channel (1.7280 ft) and non-channel cells (0.5680 ft), highlighting its accuracy and reliability. The cell-wise approach of MaxFloodCast V2, which tailors the model to the specific characteristics of each cell, allows it to better capture local variations in precipitation and terrain features, resulting in more accurate predictions. Also, the incorporation of heavy precipitation ratio features enhances its ability to capture the influence of runoff from upstream areas on channel cell depths. The dataset used for training, generated by the physics-based model, is both limited and imbalanced, comprising 90 traing events. This limitation affects the universal models more significantly, as they are less adept at handling such imbalances compared to the specialized cell-wise approach of MaxFloodCast V2. Overall, while the cell-wise model demands higher computational resources and longer processing times than universal models, its advantages outweigh these costs in the context of this pipeline, which focuses on region-specific flood depth estimation. Given its superior performance in capturing complex hydrological dynamics, the computational overhead is a minor trade-off, making MaxFloodCast V2 the optimal depth estimator for the Precipitation-Flood Depth Generative Pipeline.
@@
Figure 3: (I) Grid search results for CTGAN hyperparameters. The optimal hyperparameter set was selected based on the highest average marginal distribution scores across all synthetic precipitation-based features generated by the best-performing CTGAN configurations. The key hyperparameters explored in the grid search include the generator and discriminator learning rates, as well as the number of training epochs. (II) Stacked Marginal Distribution Comparison of Three Synthetic Features between training data and synthetic data: (a) cumulative precipitation, (b) peak precipitation and (c) duration. The distribution of the training dataset is represented in gray, while the synthetic distribution is depicted in light blue. Distributions were generated using the Synthetic Data Vault (SDV) 21.
@@ 3.2 CTGAN-generated Precipitation Records Quality Review
Among the initial 592 precipitation-flood events, we strategically sampled 90 events following the 2:4:1 ratio described in Section 2.2 and configured 36 hyperparameter sets for our grid search. This involved combining six pairs of learning rate settings with six distinct stoppage epochs, increasing from 50 epochs to 300 epochs in increments of 50, and modeling on NVIDIA’s RTX A6000 GPU. This meticulous grid search focused on generating records under the condition $\text{duration}\geq 1$, targeting the potential simulation of global precipitation events. As shown in Fig. 3I, we identified the optimal checkpoint with a generator learning rate of $10^{-2}$, a discriminator learning rate of $10^{-3}$, and early stopping at 250 epochs. Notably, the generator learning rate is set to be 10 times higher than that of the discriminator to prevent the generator from being overpowered in the early stages before it can effectively learn the underlying data patterns. Ultimately, both the generator and discriminator converged before reaching the 300-epoch limit. The optimal settings enabled a direct comparison of cumulative precipitation, peak precipitation, and duration between the training dataset and the 10,000,000 synthetic data points generated by CTGAN, as depicted in Fig. 3II. This configuration produced an average marginal distribution score of 0.802, demonstrating a high fidelity of the synthetic data, particularly for cumulative and peak precipitation when compared to the original dataset. As shown in Fig. 3II(c), we enforced a constraint in CTGAN that limits $\text{duration}\geq 1$, since our focus is primarily on global precipitation event generation. Additionally, the CTGAN model maintains the flexibility to generate local precipitation events by allowing cumulative precipitation to drop to zero. Moving forward, the challenge remains to construct multiple precipitation-depth events from the synthetic point dataset that closely correlate with training events.